Please Help Me Solve : Write A Quadratic Function In Standard Form Of (-1,0) , (0,-5) , (2,3)
Answer by scott8148(6628)

the general equation is ___ y = ax^2 + bx + c

substituting the given points will form a system of equations that can be solved for the a, b, and c coefficients

(-1,0) ___ (0) = a(-1)^2 + b(-1) + c

(0,-5) ___ (-5) = a(0)^2 + b(0) + c

(2,3) ___ (3) = a(2)^2 + b(2) + c